Find the product of 3 2/7 × 4 ⅝
Answer by Theo(13342)

find the product of 3 2/7 * 4 5/8

convert both fractions to improper decimals.

3 2/7 = 23/7

4 5/8 = 37/8

multiply 23/7 * 37/8

this is the same as (23*37) / (7*8)

result of the multiplication is 851 / 56

now you want to turn this back into a mixed fraction,.

divide 851 by 56 to get 15 with a remainder of 11.

your answer is 15 11/56